# Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning

### v1

@@ 6 Experiments
We evaluate the BiMI reward function using baseline agents (Pixl2R and ELLM-) and their BiMI-enhanced counterparts, while also exploring potential synergies with intrinsic reward models. We use the score metric for evaluation.
Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
Pixl2R $\star$ $0.142\pm 0.003$ $12.422\pm 2.439$ $9.409\pm 0.022$
Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.236\pm 2.040$ $151\%$ $10.735\pm 0.784$ $14\%$
Pixl2R + BiMI $0.162\pm 0.022$ $14\%$ $37.507\pm 7.832$ $199\%$ $7.951\pm 0.351$ $-15\%$
Pixl2R + DEIR $0.176\pm 0.009$ $23\%$ $17.372\pm 0.514$ $39\%$ $10.440\pm 1.015$ $10\%$
Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $88\%$ $57.759\pm 2.157$ $364\%$ $11.014\pm 0.190$ $17\%$
ELLM- $\star$ $0.150\pm 0.004$ $19.406\pm 10.067$ $10.826\pm 1.017$
ELLM- + Bi $0.151\pm 0.016$ $0.6\%$ $29.788\pm 1.290$ $53\%$ $11.175\pm 0.601$ $3.2\%$
ELLM- + BiMI $0.156\pm 0.014$ $4.0\%$ $33.683\pm 3.990$ $74\%$ $9.425\pm 0.267$ $-12\%$
ELLM + DEIR $0.178\pm 0.029$ $20\%$ $30.985\pm 3.507$ $59\%$ $11.857\pm 1.152$ $9.5\%$
ELLM- + BiMI + DEIR $0.279\pm 0.078$ $86\%$ $56.281\pm 6.193$ $190\%$ $13.170\pm 0.393$ $22\%$
6.1 Montezuma’s Revenge
@@
In Montezuma, Pixl2R$+$BiMI demonstrated an 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations. We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards. While this strategy effectively reduces false positives, it does not substantially mitigate the inherent reward sparsity issue in Montezuma. However, we discovered a remarkable synergy between BiMI and intrinsic reward models. While previous models showed no significant improvements with DEIR (the intrinsic reward model) alone, combining BiMI and DEIR led to a 65% performance gain. The gap in collaboration effectiveness can be attributed to two factors. In the previous setup, the consistent presence of false positive rewards misled agents towards unacceptable behaviors and hindered further exploration. Now, BiMI’s less frequent but more meaningful rewards provide anchor points for the agent’s learning. Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
@@
As illustrated in Figure 6, BiMI rewards are now concentrated on key locations. A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of false positives in the original model. Figure 9 demonstrates higher success rate grabbing the key in the first room, one of the most difficult task in Montezuma, highlighting the effectiveness of the proposed reward function and its collaboration with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks.
6.2 Minigrid
ELLM-$+$BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid. The abstract, shape-based visuals in Minigrid diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge. Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals. The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge. This capability is crucial for deploying instruction-guided agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
6.3 Crafter
We observed an intriguing pattern of results. The Bi component alone led to 14% and 3.2% improvement in performance over the original models. However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement. This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival. The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions. Furthermore, note that instruction-guided RL agents, regardless of the reward model employed, were unable to outperform pure RL agents in Crafter. This discrepancy is due to the open-world nature of Crafter, which requires dynamic strategies and real-time decision-making that our testing setups did not fully capture. Despite these challenges, it is noteworthy that Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing false positives is still beneficial across all testing environments. The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
6.4 Ablation Study and Overall Performance
@@
Our ablation study highlights the distinct contributions of the binary reward (Bi) and Mutual Information (MI) components within the BiMI framework. The binary reward mechanism alone accounted for a substantial 36.5% improvement in performance. When excluding the results from Crafter, MI component further contributes a 23% improvement over the binary reward alone. The overall improvements were substantial. BiMI led to a 67% improvement for Pixl2R and a 22% improvement for ELLM-. These results are illustrated in the line plots in Figure 8 and Figure 8 and summarized in Table 2. See Appendix A.6 for more details.
@@ 7 Conclusion
Our research reveals two key findings in instruction-guided reinforcement learning using VLM-based reward models: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function significantly mitigates the impact of noise in reward models. Our work addresses a crucial gap in prior work, paving the way for more reliable and effective instruction-guided RL systems in practical settings.
Limitation and Future Work. Our study primarily focused on linear sequences of language instructions, excluding more complex cases. Future research should investigate conditional and ambiguous instructions, which likely introduce additional challenges for VLM-based reward models. We also did not explore finetuning the VLM during agent training, a useful strategy as discussed by Fu et al. (2024).
@@ Appendix A Technical Appendix
Continued from the main text of Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning, the technical appendix consist of the following:
  • §A.1 Property of Sparse Reward Function, which discuss why sparse reward function results in the categorization of policies into a set of acceptable policies and a set of unacceptable policies but not others.
  • §A.2 Proof of the Reduction of Convergence Rate, which is referred in Section 3.
@@
  • §A.6 Additional Details of the Experiments of BiMI Reward Function, which is referred in Section 6.
  • §A.7 Cosine Similarity Score Evaluation for Different Environments, are details of the cosine similarity score evaluation for different environments, which is referred in Section 4.3.
  • §A.8 Line Plots over Average Reward for Different Environments and Ablation Study, which provides performance comparison of different models using average reward metric, the conventional metric used in RL, though they are less informative than the score metric used in the main text when dealing with sparse reward environments.
A.1 Sparse Reward Function and Range-SOAP
Formally, Abel et al. (2021) defined that:
Theorem A.1 (Abel et al. (2021)).
Proposition A.2.
Justification:
A.2 Proof of the Reduction of Convergence Rate
Specifically, the update rule of Actor-Critic algorithm is:
  • Critic:
    $\phi\leftarrow\phi-\alpha_{\phi}\nabla_{\phi}(\delta)^{2}$ (4)
@@ A.2 Proof of the Reduction of Convergence Rate
  • Actor:

### v4

@@ 7 Experiments
We continue to evaluate using Markovian Pixl2R and Non-Markovian ELLM- and their BiMI-enhanced counterparts, while also exploring potential synergies with the intrinsic reward model DEIR. We follow the same experimental setup as in Section 5.1, with additional details in Appendix A.9.
7.1 Main Results
@@
In Montezuma, Pixl2R$+$BiMI demonstrated 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations. We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards. While this strategy effectively reduces false positives, it does not substantially mitigate the inherent reward sparsity issue in Montezuma. However, we discovered a remarkable synergy between BiMI and intrinsic reward models. While previous models showed no significant improvements with DEIR (the intrinsic reward model) alone, combining BiMI and DEIR led to a 65% performance gain. The gap in collaboration effectiveness can be attributed to two factors. In the previous setup, the consistent presence of false positive rewards misled agents towards unacceptable behaviors and hindered further exploration. Now, BiMI’s less frequent but more meaningful rewards provide anchor points for the agent’s learning. Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
@@
See Figure 6 for a quantitative analysis: BiMI rewards are now concentrated on key locations. A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of false positives in the original model. Figure 7 demonstrates a higher success rate in grabbing the key in the first room, one of the most difficult tasks in Montezuma, highlighting the effectiveness of the proposed reward function and its synergy with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks. See results for other environments in Appendix A.10.
7.2 Overall Performance and Ablation Study
@@
The overall improvements were substantial. As shown in Table 2, BiMI led to a 67% improvement for Pixl2R (Markovian) and a 22% improvement for ELLM- (Non-Markovian). These results are also illustrated in Figure 17 in Appendix A.10. Our ablation study highlights the distinct contributions of the binary reward (Bi) and Mutual Information (MI) components within the BiMI framework. The binary reward mechanism alone accounted for a substantial 36.5% improvement in performance. When excluding the results from Crafter, MI component further contributes a 23% improvement over the binary reward alone.
@@ 8 Conclusion
We reveal two findings in VLM-based reward functions for RL agents: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function advocates for pessimistic rewarding, significantly mitigating the slowdown in learning caused by false positives. Our results are supported by both theoretical analysis and empirical validation across three challenging embodied tasks.
9 Limitations
Our study primarily focused on linear sequences of language instructions, excluding more complex cases. Future research should investigate conditional and ambiguous instructions, which likely introduce additional challenges for VLM-based reward models.
We also did not explore finetuning the VLM during agent training, a useful strategy as discussed by Fu et al. (2024). However, we believe that finetuning large VLMs simultaneously during reinforcement learning (RL) is computationally expensive and may not be practical in real-world scenarios.
We acknowledge that mutual information (MI) has been previously explored as a reward objective in RL. However, our work uncovers a novel insight: MI is particularly well-suited for preventing over-reliance on excessively frequent reward sources, which in turn helps to mitigate the issue of too many false positive rewards. This research is innovative in explicitly addressing the practical challenges posed by false positive rewards in instruction-following learning agents that utilize VLM rewards. We believe that our proposed BiMI reward function represents a novel application of MI specifically tailored for VLM-based rewards.
There is a gap in providing a rigorous theoretical foundation for why our theoretical findings extend to non-Markovian reward models like ELLM-. Our Assumption 4.3 facilitates the integration of non-Markovian reward models into the HuRL framework, though some might argue this assumption is strong. However, with advancements in deep RL, the distinction between non-Markovian and Markovian models has become increasingly blurred. For instance, the pioneering DQN (Mnih, 2013) utilized sequences of past observations as input, ostensibly making it non-Markovian. Yet, with deep learning’s capacity to encapsulate complex state histories into a representation that can be treated as Markovian at a higher abstraction level, some have argued that non-Markovian elements can be effectively reconsidered as Markovian within this new context (Hausknecht and Stone, 2015).
Our experiments were conducted in extremely sparse reward settings, with tasks intentionally designed to be challenging by requiring long-horizon planning and sequential execution. This setup may overemphasize the impact of false positive rewards. We acknowledge that in certain cases, rewarding partially correct trajectories may be beneficial for exploration. However, to prevent the agent from overvaluing these partial solutions, it is essential that the reward signal remains conservative, providing a pessimistic estimate of the value of partially correct trajectories.
@@ Appendix A Technical Appendix
Continued from the main text of Understanding and Mitigating Noise in VLM Rewards, the technical appendix consists of the following:
  • §A.1 Detailed Problem Setting, which provides a detailed problem setting of the instruction-following reinforcement learning. This is referred to in Section 3 and Section 4.
  • §A.2 Theoretical Analysis and Proofs, which is referred to in Section 4
  • §A.3 Additional Notes on Theoretical Setting
@@
  • §A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores, which provides extra figures that cannot be in the main body due to page limit.
  • §A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma, which provides extra evaluation results to further illustrate false positive rewards hinders agent learning.
@@
  • §A.9 Additional Implementation Details of the Experiments of BiMI Reward Function, which is referred in Section 7.
  • §A.10 Detailed Experiment Results of BiMI Reward Function, which provides detailed experiment results of BiMI reward function, which is referred in Section 7.
  • §A.11 Deprecated Convergence Analysis, It is retained here solely for reference and backtracking purposes. As the more rigorous theoretical analysis is constructed, the content of this section is now deprecated.
A.1 Detailed Problem Setting
A.1.1 The Pointer Mechanism in Reward Machine
In our problem setting, $r^{v}(\tau_{t},l_{m(t)})$ is the VLM based reward at time step $t$ evaluated using the sub-trajectory $\tau_{t}$ is the sub-trajectory up to time $t$.
Existing VLM-reward implementations applied a pointer mechanism that decide which instruction should the agent consider at the current step. We follow this implementation and denote the pointer as $m(t)$. It indicates the current instruction of the agent is trying to complete at time step $t$. $m(t)$ is updated according to the following rule:
$m(t+1)=\begin{cases}1&\text{if $t=0$}\\ m(t)+1&\text{if instr. $l_{m(t)}$ completed at $t$}\\ m(t)&\text{otherwise}\end{cases}$ (2)
In both the ELLM and the Pixl2R framework that uses VLM-based rewards, the assessment of whether instruction $l_{m(t)}$ is completed at time $t$ is based on the cumulative reward for that instruction. The pointer remains on the current instruction until the accumulated reward reaches a predetermined threshold. For a detailed pseudo-code of the existing VLM reward + RL algorithm, please refer to Appendix A.4.2. Note that our proposed BiMI reward function uses a different mechanism to determine if instruction $l_{m(t)}$ is completed at time $t$.
In logic, an atomic sentence is a type of statement which cannot be broken down into other simpler sentences. While the pointer mechanism effectively enforces the sequential ordering of instructions at a higher level, each individual instruction within the expert walkthrough is not necessarily atomic. This means that a single instruction can encapsulate multiple finer-grained requirements for the agent’s behavior. Therefore, even though the pointer mechanism is implemented, the internal composition of each instruction is not strictly enforced. Consequently, the issue of “composition insensitivity” can arise when the VLM tries to align agents’ trajectories with non-atomic instruction sentences.
Just as potential-based reward shaping (PBRS) modifies the original MDP into a reshaped MDP through the use of auxiliary rewards, the HuRL framework similarly transforms the original MDP $\mathcal{M}$ into a new MDP ~$\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\widetilde{r},\widetilde{\gamma}\rangle$, employing a heuristic-based auxiliary reward $h$, along with a coefficient $\beta\in[0,1]$ that scales the auxiliary rewards. Thus, the updated reward function can be written as:
$\widetilde{r}(s,a)=r^{e}(s,a,s^{\prime})+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a}[h(s^{\prime})]\quad\text{and $\widetilde{\gamma}=\beta\gamma$}$
We also provide the formal definition of false negative reward as below.
Definition A.1 (False Negative Rewards).
A false negative reward occurs when:
Instruction-Following Perspective: The VLM-based reward $r^{v}(\tau_{t},l_{m(t)})\approx 0$ for a trajectory $\tau_{t}$ that does satisfy instruction $l_{m(t)}$.
Heuristic Perspective: The heuristic $h(s_{t})\approx 0<V^{*}(s_{t})$, underestimating the optimal value of $s_{t}$.
A.1.2 Detailed Terminology
This section contains detailed terminology from both the standard RL and the HuRL framework.
We start by setting clear the terminology. The original MDP is defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},r^{e},\gamma\rangle$ as set out in the main text. Let $G^{e}(\tau)$ denote the cumulative environment reward of a single trajectory $\tau=\langle s_{0},a_{0},\ldots,s_{T}\rangle$ where $T$ is the total time of the trajectory, we have
$G^{e}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{e}(s_{t},a_{t}).$ (3)
A policy $\pi_{\theta}$ is a conditional distribution $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ parameterized by $\theta$, where $\Delta(\cdot)$ denotes the space of probability distribution. The probability of a trajectory $\tau$ being generated, given that actions come from $\pi_{\theta}$ and the starting state is $s_{0}$, is expressed as:
$P(\tau\mid\pi_{\theta})=\prod_{t=0}^{T}\mathcal{P}(s_{t+1}\mid s_{t},a_{t})\pi_{\theta}(a_{t}\mid s_{t})$ (4)
Because a policy $\pi$ is stochastic, different trajectories could be generated from the same policy for a given initial state. The objective function to be optimized is the expected cumulative environment reward of trajectories generated by $\pi$ starting from state $s_{0}$. It is also called the state value function of $s_{0}$ and denoted as:
$V^{\pi}(s_{0}):=\mathbb{E}_{\rho_{s}^{\pi}}\left[G^{e}(\tau)\right]=\sum_{\textnormal{all $\tau_{i}$ generated by $\pi_{\theta}$}}G^{e}(\tau_{i})\cdot P(\tau_{i}|\pi)$ (5)
where $\rho_{s}^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}=s$. We defined an acceptable policy $\pi_{A}$ as
$\pi_{A}:V^{\pi_{A}}\geq V^{\pi^{*}}-\epsilon$ (6)
where $\pi^{*}$ is the optimal policy and $\epsilon$ is a control parameter that quantifies the acceptance of other policies. Acceptable policies defined in this way are also referred to as $\epsilon$-optimal policies. We set $V^{*}=V^{\pi^{*}}$ for easy reference. Any policy that is not an acceptable policy is an unacceptable policy $\pi_{U}$. We define a good trajectory $\tau_{G}$ as
$\tau_{G}:G^{e}(\tau_{G})\geq V^{*}-\epsilon.$ (7)
Any trajectory that is not a good trajectory is a bad trajectory $\tau_{B}$. Note that this way of distinguishing good/bad trajectories and acceptable/unacceptable policies is explicitly based on the environment reward $r^{e}$, and introducing additional or alternative rewards into the system does not change this definition. For easy reference, we call any trajectory that reaches a goal state a goal trajectory, and we call any policy that can generate at least one goal trajectory a goal policy.
We denote the state distribution of a policy $\pi$ at time $t$ as $d^{\pi}_{t}$. Thus, the discounted average state distribution of a policy $\pi$ can be expressed as $d^{\pi}=(1-\gamma)\sum^{\infty}_{t=0}\gamma^{t}d^{\pi}_{t}$, where $(1-\gamma)$ is the normalization factor to ensure the result is a proper probability distribution. We also define a shorthand notation – for a state distribution $d\in\Delta(\mathcal{S})$, we define $V(d)=\mathbb{E}_{s\sim d}\left[V(s)\right]$.
When the Heuristic-Guided Reinforcement Learning (HuRL) framework is applied, the reshaped MDP $\widetilde{\mathcal{M}}$ has separated reshaped reward $\widetilde{r}$ and separated discount factor $\widetilde{\gamma}=\beta\gamma$, where $\beta\in[0,1]$ is the hyperparameter that scales the auxiliary heuristic rewards. We denote the value function of the reshaped MDP as $\widetilde{V}$, and the optimal policy and its value function under $\widetilde{\mathcal{M}}$ as $\widetilde{\pi}^{*}$ and $\widetilde{V}^{*}(\widetilde{V}^{\widetilde{\pi}^{*}})$.
The discussion regarding the convergence guarantee of learning $\pi$ in $\widetilde{\mathcal{M}}$ resulting in the optimal policy in the original MDP $\mathcal{M}$ can be found in Appendix A.3.1. For the purposes of this paper, it is sufficient to note that the authors of HuRL have provided a trick to ensure convergence. Moreover, empirical experiments conducted by the authors demonstrate that HuRL converges even without this trick.
It is also important to distinguish the Bellman backup equation under the two different MDPs: the original $\mathcal{M}$ and the reshaped $\widetilde{\mathcal{M}}$. By definition, $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. In contrast, (~$(\widetilde{\mathcal{B}}h)(s,a)=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. Nevertheless, the two Bellman backup equations possess a remarkable property – they are essentially equivalent to each other:
(~$(\widetilde{\mathcal{B}}h)(s,a)$ $=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
$=\left(r(s,a)+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]\right)+\beta\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
$=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
$=(\mathcal{B}h)(s,a)$ (8)
A.1.3 Detailed Construction of Markovian Heuristic using Non-Markovian Rewards from VLM
Specifically,
$h(s_{t}):=\begin{cases}A(s_{t})V^{*}(s_{t})&\text{if $s_{t}\in S^{m(t)}$}\\ 0&\text{otherwise}\end{cases}$ (9)
@@ A.1.3 Detailed Construction of Markovian Heuristic using Non-Markovian Rewards from VLM
$A(s_{t}):=\frac{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau_{t^{\prime}},l_{m(t^{\prime})})}{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau^{*}_{t^{\prime}},l_{m(t^{\prime})})}\in[0,1]$ (10)
is a scaling factor that quantifies how closely the agent is following previous instructions.
A.2 Theoretical Analysis and Proofs
A.2.1 Convergence time on a sparse-reward landscape
We can pin down a characteristic convergence time on a sparse-reward landscape. The sparse-reward setting enforces that
$r^{e}(s_{t})=\begin{cases}1&\text{if $s_{t}\in\mathcal{S}_{G}$}\\ 0&\text{otherwise}\end{cases}.$ (11)
It is clear from the definition that a good trajectory must always be a goal trajectory, and the cumulative reward is simply the reward at the final goal state so $G^{e}(\tau)=\gamma^{T}$ in this setting. For a randomly initialized policy, it is highly unlikely that the initial distribution of trajectories contains any goal trajectory due to the sparsity of goal states. The optimization of $V^{e}_{\pi_{\theta}}$ thus consists of two parts. The first part is to search for a goal trajectory. The gradient landscape is almost 0 everywhere, except for cases where a trajectory is $\delta$-close to a goal trajectory. Here $\delta$ is the differential unit in the numerical differentiation used in the gradient calculation
$\theta=\theta+\alpha\nabla_{\theta}V_{\pi_{\theta}}$ (12)
such that $\delta$-close means being numerically accessible within a distance of $|\delta|$ in parameter space. And the second part is to reduce $T$ so that goal trajectories become good trajectories and consequently achieving acceptable policies. For the first part, searching for a trajectory for the target is effectively a random walk in the $d$-dimensional parameter space due to the flat gradient landscape.
Lemma A.2.
For a random walk in $n$-dimensional space, the expected number of steps $T_{D}$ needed to travel a distance of $D$ scales with $D^{2}$.
Proof: let $\vec{X_{1}}$, $\vec{X_{2}}$, … $\vec{X_{T}}$ be IID random unit vectors uniformly distributed on a $(d-1)$-dimensional sphere $\mathcal{S}^{d-1}\subset\mathbb{R}^{d}$, where $\vec{X_{i}}=(X_{i1},...,X_{id})$ and $|\vec{X_{i}}|^{2}=\sum_{j=1}^{d}X^{2}_{ij}=1$. Let $\vec{S_{T}}:=\sum_{i=1}^{T}\vec{X_{i}}$. By a $n$-dimensional cosine rule we have
$|\vec{S}_{T}|^{2}$ $=|\vec{S}_{T-1}|^{2}+2\vec{S}_{T-1}\cdot\vec{X}_{T}+|\vec{X}_{T}|^{2},$ (13)
and because $\mathbb{E}[\vec{X}_{T}]=\vec{0}$
$\mathbb{E}[|\vec{S}_{T}|^{2}]$ $=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+\mathbb{E}[2\vec{S}_{T-1}\cdot\vec{X}_{T}]+1$ (14)
$=\mathbb{E}[|\vec{S}_{T-1}|^{2})]+2\vec{S}_{T-1}\mathbb{E}[\vec{X}_{T}]+1$ (15)
$=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+1$ (16)
$=\mathbb{E}[|\vec{S}_{T-2}|^{2}]+1+1$ (17)
$\vdots$ (18)
$=T$ (19)
i.e. $\mathbb{E}[|\vec{S}_{T}|]\sim\sqrt{T}$. When a policy is randomly initialized with $\theta_{0}$, the distance $D$ to a goal policy, $D:=||\theta_{goal}-\theta_{0}||$ is fixed and is the distance the random walk needs to travel ($\mathbb{E}[|\vec{S}_{T}|]=D\sim\sqrt{T_{D}}$), so the characteristic time needed to travel this distance $T_{D}\sim D^{2}$ as we have shown above.
See 4.1 Auxiliary rewards essentially open the path for a divide-and-conquer approach by introducing intermediate rewards in the learning process. We introduce BiMI rewards as
and the cumulative reward of a single trajectory in the presence of BiMI rewards becomes
$G^{v}_{\textsc{BiMI}}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{v}_{\textsc{BiMI}}(\tau,l_{m(t)}).$ (21)
[…]
See 4.2 Proof: the expected time taken for each of the sub-tasks then scales with ${}d_{i}^{2}$ respectively (Lemma A.2), and we have
$d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2}$ $<(d_{1}+d_{2}+...+d_{n-1})^{2}$ (22)
$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]$ $<\mathbb{E}[T_{D}]$ (23)
because $d_{i}>0\forall i$. We can also work out that the upper bound for this improvement is a factor of $n-1$ by invoking the Cauchy–Schwarz inequality $\left(\sum_{i=1}^{n}u_{i}^{2}\right)\left(\sum_{i=1}^{n}v_{i}^{2}\right)\geq\left(\sum_{i=1}^{n}u_{i}v_{i}\right)^{2}$:
$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})(1^{2}+1^{2}+...+1^{2})$ $\geq(d_{1}+d_{2}+...+d_{n-1})^{2}$ (24)
$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})$ $\geq\frac{1}{n-1}D^{2}$ (25)
$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]\geq\frac{1}{n-1}\mathbb{E}[T_{D}]$ (26)
and the equality sign holds (indicating maximal improvement) when $d_{1}=d_{2}=...=d_{n-1}$. The intuitive interpretation is that the divide-and-conquer approach is the most effective when the task is divided evenly into subtasks. Additionally, the presence of $p(l_{k})$ introduces some lexical-level tolerance in the learning process and results in a considerably larger $\delta$-close radius which could result in further improvement in the convergence rate. However, as we will discuss in the next section, introducing too much tolerance could harm the convergence rate rather than improve it.
A.2.2 False Positive and the Violation of Pessimistic Property of Heuristics
See 4.7
Proof.
We want to show that given: